# Crossing the Human-Robot Embodiment Gap with Sim-to-Real RL using One Human Demonstration

### v2

@@ Abstract
Teaching robots dexterous manipulation skills often requires collecting hundreds of demonstrations using wearables or teleoperation, a process that is challenging to scale. Videos of human-object interactions are easier to collect and scale, but leveraging them directly for robot learning is difficult due to the lack of explicit action labels from videos and morphological differences between robot and human hands. We propose Human2Sim2Robot, a novel real-to-sim-to-real framework for training dexterous manipulation policies using only one RGB-D video of a human demonstrating a task. Our method utilizes reinforcement learning (RL) in simulation to cross the human-robot embodiment gap without relying on wearables, teleoperation, or large-scale data collection typically necessary for imitation learning methods. From the demonstration, we extract two task-specific components: (1) the object pose trajectory to define an object-centric, embodiment-agnostic reward function, and (2) the pre-manipulation hand pose to initialize and guide exploration during RL training. We found that these two components are highly effective for learning the desired task, eliminating the need for task-specific reward shaping and tuning. We demonstrate that Human2Sim2Robot outperforms object-aware open-loop trajectory replay by 55% and imitation learning with data augmentation by 68% across grasping, non-prehensile manipulation, and multi-step tasks. Project Site: human2sim2robot.github.io
I Introduction
Training robotic manipulation policies from demonstrations via imitation learning (IL) has gathered considerable interest [1, 2, 12, 19, 23]. A key requirement for effective IL is the availability of high-quality demonstrations composed of sequences of states with corresponding robot action labels. This is typically acquired through labor-intensive data collection requiring hundreds or thousands of human demonstrations per task via teleoperation or specialized wearable equipment [10, 28, 53]. In contrast, videos of humans interacting with objects with their own hands are inexpensive to collect, a scalable alternative to traditional demonstration collection methods. However, leveraging human videos directly for IL is challenging as they lack explicit robot action labels [57].
@@ I Introduction
Figure 1: Our Framework. Human2Sim2Robot is a real-to-sim-to-real RL framework for learning dexterous manipulation policies from one human RGB-D video demonstration per task. From the demonstration, we leverage object pose trajectories and pre-manipulation hand poses to provide guidance and advantageous initialization for RL training in simulation. These policies demonstrate zero-shot sim-to-real transfer to a Kuka arm and Allegro hand.
One popular approach to address this is by obtaining per-timestep human hand pose estimates and converting them to robot action labels via fingertip retargeting and inverse kinematics (IK) [4, 53]. However, this approach faces two challenges. First, while recent hand pose reconstruction methods have shown impressive performance [38], their predictions are still not always accurate due to occlusions and sensor noise, resulting in mismatches between the robot hand IK targets and the true human hand fingertips. Second, even with perfect human hand pose estimates, the considerable embodiment gap between human and robot hands may make solving IK infeasible or result in suboptimal robot joint configurations for the demonstrated task. Both challenges are particularly punishing for contact-rich, dexterous manipulation tasks using multi-fingered robotic hands as they require precise coordination of forces and motion at multiple contact points, which is difficult to achieve when retargeted actions are noisy or suboptimal for the target embodiment [3, 32, 37]. Existing IL methods are ill-equipped to address this issue as they rely on accurate correspondences between the demonstrated and learned behaviors. Reinforcement learning (RL) offers a promising alternative to overcome these limitations by enabling the robot to directly learn manipulation tasks using its own embodiment. However, RL has several limitations such as tedious, task-specific reward engineering and unfavorable sample complexity, which make training real-world policies infeasible [49, 60].
@@
We propose Human2Sim2Robot, a real-to-sim-to-real RL framework that addresses these limitations by learning robust dexterous manipulation policies from a single human hand RGB-D video demonstration. We extract two task-specific components from the video: (1) the object 6D pose trajectory, which defines an object-centric reward function that specifies the desired task while accommodating embodiment differences, and (2) the pre-manipulation hand pose, which provides advantageous initialization of RL training such that the resulting policy converges on a similar manipulation strategy shown in the demonstration. Formalizing an expert task demonstration with these two components facilitates RL policy training with no task-specific reward tuning. Instead of directly learning state-action mappings, we use the demonstration for task specification and guidance, encouraging human-like behavior when beneficial while allowing deviations when the human strategy is unsuitable for the robot’s embodiment. Human2Sim2Robot policies achieve zero-shot sim-to-real transfer on a real dexterous robot, without requiring wearable technologies, teleoperation, or large-scale data collection.
@@
To the best of our knowledge, Human2Sim2Robot is the first system that learns a robust real-world dexterous manipulation policy from only one human RGB-D video demonstration, bridging the human-robot embodiment gap for a wide range of grasping, non-prehensile manipulation, and multi-step tasks. Our extensive ablation studies demonstrate the importance of our system’s design decisions. While individual components have precedents, these works are often limited to simulation [42, 47], are not reactive closed-loop policies [6, 22, 56], require significantly more demonstrations [25, 42, 47, 58], or only perform prehensile manipulation [42, 47, 56, 22]. We demonstrate that Human2Sim2Robot is highly effective for training task-specific RL policies in simulation. The resulting policies can execute diverse real-world dexterous manipulation tasks, such as pouring from a pitcher, pivoting a box against a wall, inserting a plate into a dishrack, or even composed sequences of tasks, without any task-specific reward tuning. In the challenging single demonstration regime, Human2Sim2Robot outperforms object-aware open-loop trajectory replay by 55% and imitation learning with data augmentation by 68% across grasping, non-prehensile and extrinsic manipulation, and complex multi-step tasks.
II Related Work
II-A Imitation Learning for Robotics
II-A1 Visuomotor Imitation Learning

### v3

@@ Abstract
Teaching robots dexterous manipulation skills often requires collecting hundreds of demonstrations using wearables or teleoperation, a process that is challenging to scale. Videos of human-object interactions are easier to collect and scale, but leveraging them directly for robot learning is difficult due to the lack of explicit action labels and human-robot embodiment differences. We propose Human2Sim2Robot, a novel real-to-sim-to-real framework for training dexterous manipulation policies using only one RGB-D video of a human demonstrating a task. Our method utilizes reinforcement learning (RL) in simulation to cross the embodiment gap without relying on wearables, teleoperation, or large-scale data collection. From the video, we extract: (1) the object pose trajectory to define an object-centric, embodiment-agnostic reward, and (2) the pre-manipulation hand pose to initialize and guide exploration during RL training. These components enable effective policy learning without any task-specific reward tuning. In the single human demo regime, Human2Sim2Robot outperforms object-aware replay by over 55% and imitation learning by over 68% on grasping, non-prehensile manipulation, and multi-step tasks. Website: human2sim2robot.github.io
Keywords: Dexterous Manipulation, Reinforcement Learning, Sim-to-Real
1 Introduction
@@
Figure 1: Our Framework. Human2Sim2Robot learns dexterous manipulation policies from one human RGB-D video using object pose trajectories and pre-manipulation poses. These policies are trained with RL in simulation and transfer zero-shot to a real robot.
Human-like dexterous hands have the potential to significantly advance robotic manipulation [1, 2, 3]. However, the complexity of dexterous robot hands introduces substantial challenges for many existing robot learning methods. For example, imitation learning (IL) from human demonstration has shown success using simpler end-effectors with large amounts of training data [4, 5, 6, 7, 8], but collecting high-quality demonstrations for dexterous hands is far more difficult. Capturing high-quality 3D human hand motion typically relies on wearable sensors and teleoperation systems [9, 10], which are expensive and difficult to scale.
In contrast, videos of humans interacting with objects using their own hands are inexpensive to collect and offer a scalable alternative to traditional demonstration collection. However, leveraging them directly for robotic IL is challenging as they lack explicit robot action labels [11]. One common approach to address this is by obtaining per-timestep human hand pose estimates and converting them to robot action labels via fingertip retargeting and inverse kinematics (IK) [9, 10]. However, this approach is often unreliable as hand pose reconstruction methods [12] are susceptible to occlusion and sensor noise. Even with perfect pose estimates, this simple retargeting strategy often results in suboptimal robot trajectories due to morphological differences between the robot and human. These challenges are particularly punishing for contact-rich, dexterous manipulation [13, 14, 15].
Existing IL methods are ill-equipped to address this issue as they rely on accurate correspondences between demonstrated and learned behaviors. Reinforcement learning (RL) offers a promising alternative to overcome these limitations by enabling robots to directly learn manipulation tasks using their own embodiment. However, RL has several limitations such as tedious, task-specific reward engineering and unfavorable sample complexity, making real-world policy training infeasible [16, 17].
In this paper, we propose Human2Sim2Robot, a real-to-sim-to-real RL framework that addresses the limitations of existing methods and combines the best of both worlds: it only requires a single human RGB-D video demonstration and does not require any task-specific reward engineering. Crucially, we found that high-fidelity 3D human motion data is not necessary to learn robust dexterous manipulation policies. Instead, training an RL dexterous manipulation policy only requires two task-specific components that can be reliably extracted from the human video: (1) the object 6D pose trajectory, and (2) a single pre-manipulation hand pose.
@@
We use (1) to define an embodiment-agnostic, object-centric reward that specifies the desired task, and (2) to provide advantageous initialization for RL training and facilitate more efficient exploration. Formalizing an expert demonstration with these two components facilitates RL policy training with no task-specific reward tuning. Instead of directly learning state-action mappings, we use the demonstration for task specification and guidance, encouraging human-like behavior while allowing deviations when the human strategy is unsuitable for the robot’s embodiment. This enables Human2Sim2Robot policies to achieve zero-shot sim-to-real transfer on a real-world dexterous robot, without requiring wearables, teleoperation, or large-scale data collection.
@@
To the best of our knowledge, Human2Sim2Robot is the first system that learns a robust real-world dexterous manipulation policy from only one human RGB-D video demonstration, bridging the human-robot embodiment gap across grasping, non-prehensile manipulation, and complex multi-step tasks. We achieve this with just a few minutes of human effort end-to-end, from demonstration collection to digital twin construction. Our extensive ablation studies demonstrate the importance of our system’s design decisions; while individual components have precedents, these works are often limited to simulation [1, 18], are not reactive closed-loop policies [3, 19, 20], require significantly more demonstrations [1, 18, 21, 22], or only perform prehensile manipulation [1, 18, 19, 20].
Human2Sim2Robot policies can execute diverse real-world dexterous manipulation tasks, such as pouring from a pitcher, pivoting a box against a wall, and inserting a plate into a dishrack, without any task-specific reward tuning. In the single human demo regime, our method outperforms object-aware trajectory replay by >>55% and imitation learning by >>68% across all real-world tasks.
2 Related Work